A handy man is installing a new shelf in his work shop. He wants to cut a 72 inch board into 3 pieces, so that each shelf would be 6 inches longer than the proceeding board. What is the length of each shelf?
Answer by Earlsdon(6294)

Let x = the length of the shortest piece, then x+6 = the length of the second piece and x+12 = the length of the longest piece. The sum of the lengths of these three pieces is 72 inches, so...
x+(x+6)+(x+12) = 72 Simplify.
3x+18 = 72 Subtract 18.
3x = 54 Divide by 3.
x = 18 inches.
x+6 = 24 inches.
x+12 = 30 inches.